# NNN: Next-Generation Neural Networks for Marketing Mix Modeling

### v2

@@
NNN: Next-Generation Neural Networks for Marketing Mix Modeling
@@ 1 Abstract
We present NNN, a Transformer-based neural network approach to Marketing Mix Modeling (MMM) designed to address key limitations of traditional methods. Unlike conventional MMMs which rely on scalar inputs and parametric decay functions, NNN uses rich embeddings to capture both quantitative and qualitative aspects of marketing and organic channels (e.g., search queries, ad creatives). This, combined with its attention mechanism, enables NNN to model complex interactions, capture long-term effects, and potentially improve sales attribution accuracy. We show that L1 regularization permits the use of such expressive models in typical data-constrained settings. Evaluating NNN on simulated and real-world data demonstrates its efficacy, particularly through considerable improvement in predictive power. Beyond attribution, NNN provides valuable, complementary insights through model probing, such as evaluating keyword or creative effectiveness, enhancing model interpretability.
@@ 2 Introduction
Marketing Mix Models (MMMs), with roots tracing back to the 1950s [19, 4], have seen a resurgence in recent years [7, 23, 13, 45], offering a means to estimate the effectiveness of diverse media channels using aggregated, observational data. This renewed interest is partly fueled by the increasing difficulty of long-term user-level tracking for attribution or experiments [32], especially given the deprecation of third-party cookies and stricter privacy regulations like GDPR, CCPA [6] and Apple’s Intelligent Tracking Prevention (ITP) [1]. MMMs circumvent these tracking needs by operating on aggregated data (e.g., weekly impressions and sales per region).
However, traditional MMMs face well-documented limitations [7, 33]. First, capturing long-term brand effects is notoriously difficult; parametric adstock functions typically used to model carryover effects often lose identifiability from the baseline beyond a few months. Second, MMM datasets are frequently sparse relative to the number of factors influencing sales, which can lead to statistically overdetermined models [7] and has historically limited the feasibility of employing more complex, data-hungry techniques like deep neural networks. Third, the common assumption [^1] There are a few exceptions such as [30]. of time-invariant response curves makes it challenging to account for variations in marketing effectiveness due to changes in ad creative quality, targeting strategies, or external factors like seasonality [^2] Such as changes in effectiveness during the holiday shopping season..
To address these challenges, this paper introduces NNN (Next-Generation Neural Networks for Mixed Media Modeling), a novel paradigm leveraging recent advancements in deep learning and representation learning. NNN departs from traditional MMMs in several key ways:
  • Rich Data Representation. Instead of relying solely on scalar inputs (e.g., spend), NNN utilizes high-dimensional embeddings to represent both marketing activities (e.g., Search Ads, YouTube Ads) and crucial organic signals (e.g., Google Search queries). These embeddings capture not only quantitative volume but also qualitative aspects like ad creative content or search query semantics, allowing the model to directly address the challenge of time-varying creative effectiveness.
@@
Our paper is structured as follows. We give an overview of related working influencing our approach (Related Work). Next, we discuss the data used (Datasets), and how they are formatted in our method (Data Structure). We give an overview of the computation of the model (Model Overview), then go into further details regarding the how traditional data is augmented and embedded (Input Data Representations). We discuss the intricacies of our Transformer (Transformer) and details about how the model makes the final predictions (Prediction Heads). We discuss training details (Model Training) and attribution methods (Sales Attribution). We give an overview of our experiments and then discuss the results and analysis (Experiments). Finally, we give a discussion about how our method fits into the existing world of MMMs (Discussion), and give a concluding overview (Conclusion).
@@ 3 Related Work
Our proposed NNN framework builds upon and diverges from several strands of research in marketing modeling, primarily traditional and Bayesian Media Mix Modeling, the application of neural networks in this domain, and the use of embeddings to represent media signals. This section situates NNN within this existing literature.
@@ 3.1 Media Mix Modeling Evolution
The field of Media Mix Modeling (MMM) provides the foundational context for NNN, offering established methods for assessing advertising ROI using historical, aggregate data [19, 4]. Recognizing limitations in early linear models, Bayesian Media Mix Modeling (BMMM) emerged, introducing techniques to model ad carryover and saturation (shape effects) more flexibly, often using Hill and adstock functions [23]. BMMM leverages Bayesian inference (e.g., MCMC methods) to incorporate prior knowledge into model estimation, a crucial aspect given the often data-constrained nature of MMM problems. This Bayesian framework was further extended to handle geographical variations through hierarchical modeling [34]. Meridian, a Google Bayesian MMM framework integrated advancements of MMM research over the past decade. It incorporates features like ROI priors and calibration [44], reach and frequency data integration [43], and paid search modeling. While powerful, these methods typically rely on scalar inputs and pre-defined functional forms for temporal effects, motivating our exploration of alternative architectures.
@@ 3.2 Neural Networks in Media Mix Modeling
The potential for neural networks (NNs) in MMM has been recognized for some time [36], owing to their ability to capture complex non-linearities. However, concerns over interpretability and the data requirements for training large NNs have historically limited their adoption in practice. Recent efforts aim to bridge this gap. For instance, CausalMMM [15] attempts to learn causal structures using Granger causality [18] and employs a variational autoencoder (VAE) [12] with a dedicated causal relational encoder and a marketing response decoder where the temporal response is learned alongside the saturation response. While innovative in its use of NNs and causality, this approach, like most traditional MMMs, still operated on scalar representations of media activity. NNN builds on the potential of NNs but focuses on leveraging richer input representations via embeddings.
@@ 3.3 Embedding-Based Media Representations
A core innovation of NNN is its use of high-dimensional embeddings, drawing heavily on the work presented in [28]. That work demonstrated that organic signals, specifically aggregated Google Search query embeddings, could be highly predictive of real-world outcomes like flu rates and auto sales. This contrasted with earlier studies that used Google Search data but often relied on counts of specific keywords, necessitating significant feature engineering [9, 26, 25, 38, 40]. The key insight from [28] was using pre-trained language models to embed search terms and then aggregating these embeddings (via summation in SLaM) to create a dense “search embedding” for a given time and geo, capturing both the volume and semantic content of search activity as a proxy for organic consumer demand. In our work, we show that the search embedding is highly effective baseline for MMM models, and we introduce additional embeddings for other paid marketing channels such as Google Search ads and YouTube ads.
@@ 3.4 Terminology and Conventions
Throughout this paper, we adopt specific terminology for clarity. We use “MMM” to refer generally to traditional Marketing Mix Models, often exemplified by frameworks like Meridian. In contrast, “NNN” denotes our proposed neural network-based methodology.
@@ 4 Datasets
    Data was randomly generated using the test_utils module of the Meridian package to simulate 130 weeks of data over 100 random geos. We created two channels of random media data (YouTube and Search Ads) and one channel of random controls data (Search), applying adstock and Hill functions to the media channels. In order to model sales driven directly by YouTube, we assumed an adstock retention rate $\alpha_{m}=0.75$ and a Hill function effective concentration $e_{c}=1.0$ and applied these functions to the random YouTube media data with arbitrary normalization. For the indirect sales driven by YouTube, we applied a different set of adstock and Hill functions to the YouTube media data ($\alpha_{m}=0.5$ and $e_{c}=3$) and added those normalized impressions to the random search data so that they contributed an additional 20% or so on top of the random search data. We then generated sales from the sum of the random search impressions and the YouTube-generated search impressions, assuming a flat conversion rate of 0.1%. Finally, search ads (which in this simplified model are not causally connected to search) also generated sales using $\alpha_{m}=0.3$ and $e_{c}=1$.
@@
    In the end we have two versions of this dataset: High variance and low variance, differing only in the magnitude of the random noise corresponding to “intent” or creative effectiveness. For the high-variance dataset, this varies uniformly between 0.5 and 1.5, while in the low-variance dataset it varies uniformly between 0.8 and 1.2. We compare our method against traditional MMMs on these datasets on both sales-prediction error-rate evaluation and on sales channel attribution estimates, because the ground-truth attributions are known when creating the data.
  • Real-world, Regional-level. (4 regions). This data is real-world marketing and sales data, aggregated at a regional- and weekly-level. We use it for an apples-to-apples comparison against traditional MMM in real-world setting, where previous MMMs have already been run, and priors have been established. We use only the marketing volume as model inputs (quantitative information) in both NNN and traditional MMM. The Search embeddings are generated using approximated 8M queries that are categorically related the business’s sales.
@@ 6 Model Overview
Our baseline NNN model is based on the Search-to-Sales dependency in Figure 1. Previous work has shown that Google Search is a good proxy for the consumer demand, and even captures seasonal effects; we use an embedded version of Search where user queries are aggregated according to their semantic representations from a language model [28]. We pass this embedded representation of search to a function that is similar to the hill functions found in traditional MMMs to get our estimate of Sales. We enrich our Sales estimate by including marketing-to-search effects (the blue lines from Marketing to Sales in 1) in a similar fashion, utilizing embedded versions of the paid media. To include the temporal effects in Figure 1 (the yellow lines from Search to Sales and the blue lines from Marketing to Sales), we utilize a Transformer architecture that is able to attend to channels’ histories. This allows the feature representation used to predict sales to depend not just on the current time-period’s Search, but also it’s history. Similarly, this is done for all other channels. Finally, our model is allowed to be trained not just to predict sales, but to predict organic channels like Search (the grey lines in Figure 1). This allows our model to capture marketing effects that happen through an intermediary such as Google Search (E.g., Imagine a channel that affects Search, and which then drives sales).
@@ 7.5 Scalars
In the case where there is no way to use a model to embed data (e.g., in the scenario where only impressions of a channel are stored, and no other contextual information is stored along side it, perhaps for privacy reasons or legacy logging methods), similar to traditional MMMs, we use only the scalar quantity in our models. Our architecture requires all channels to be represented as a $D$-dimensional value, so we pad the scalar with $D-1$ zero values for the other dimensions. Note that the magnitude of this vector is still the channel volume, which is consistent with all other channels.
@@ 8.1 Attention
In our model, we use historical information across marketing channels to predict sales or other business KPIs on a given day $t$. We employ attention mechanisms [2, 39] that learn which channels and prior time steps are most relevant for the current prediction context (channel $c$, time $t$).
@@ 9.1 Sales Head
and the function $P$ is an MLP Resnet where the final layer has one hidden unit and a sigmoid activation function with parameters $\theta^{\text{Search}}$. This structure is similar to Hill functions in that it saturates, however in typical MMMs, the saturation is with respect to the advertising volume. In this model, the volume scales linearly without saturation, and the output of the saturated function represents the conversion rate of a Search query.
To add multi-channel flexibility in our model (at first), we incorporate the effect of more marketing channels additively [^5] We experimented with using multiplicative effects via using a log-scale version of the model, but we found the additive models generally performed better and match our expectations that most marketing channels are independent of one another; see the Appendix for more details as
@@ 9.2 Search Head
It is long believed that intermediary organic channels such as Google Search are moved via marketing, but are often neglected to be measured. We aim to incorporate these effects into NNN by generalizing our modeling framework to predict not just sales, but also the Search information. We do so by predicting the search embedding $X_{g,t+1,\text{Search}}$ at time step $t$ for geo $g$, learning both the magnitude and direction of the search [^7] This means our model output has values that are representative of the following time-step for the Search channel.. This allows us to capture movements in search patterns due to marketing (i.e., movements in the “baseline” sales) which are uncaptured in traditional MMM, and have been a major point of scrutiny. Our search head looks like
@@ 9.3 Other Heads
The other inputs to our model are marketing channels, which are controlled inputs, whereas the Search and Sales channels are measurements / observations. Because we treat these as marketing levers and not observations (although it is a viewpoint that Search can affect Search ads, we neglect this in our models), we don’t have a need to predict their values during training. Out of convenience, to keep the tensor dimensions consistent, we project the intermediate representations of all other channels back to their original size.
@@ 10 Model Training
All traditional MMMs are trained using Meridian. For the simulation experiments, we use a single knot and Meridian’s default priors. For the real-world experiments we use priors from a previous training of the model, which replicates a common practice with real-world MMMs. Additionally, we use the same number of knots that were previously used to fit the model, with an additional two knots fit during the test period, which are designed to help the model generalize to the test period.
@@ 11 Sales Attribution
where $X$ represents the original data and media volume and $\tilde{X}^{c}$ represents the same data except the media from channel $c$ is set to zero. Because there can be bias in predictions, we report the predictions on a mix or percentage basis, defined as $\text{mix}\%_{c}=\frac{\text{Attribution}_{c}}{\sum_{c^{\prime}}\text{Attribution}_{c}^{\prime}}$.
@@
Since our model also allows us to model the effect on the intermediary organic channels, we can autoregressively unroll predictions out in time. We start with a prefix $P$ of data and compute the following

### v3

# NNN: Next-Generation Neural Networks for Marketing Measurement

@@
NNN: Next-Generation Neural Networks for Marketing Measurement
@@ 1 Abstract
We present NNN, an experimental Transformer-based neural network approach to marketing measurement. Unlike Marketing Mix Models (MMMs) which rely on scalar inputs and parametric decay functions, NNN uses rich embeddings to capture both quantitative and qualitative aspects of marketing and organic channels (e.g., search queries, ad creatives). This, combined with its attention mechanism, potentially enables NNN to model complex interactions, capture long-term effects, and improve sales attribution accuracy. We show that L1 regularization permits the use of such expressive models in typical data-constrained settings. Evaluating NNN on simulated and real-world data demonstrates its efficacy, particularly through considerable improvement in predictive power. In addition to marketing measurement, the NNN framework can provide valuable, complementary insights through model probing, such as evaluating keyword or creative effectiveness.
@@ 2 Introduction
Marketing Mix Models (MMMs), with roots tracing back to the 1950s [19, 4], have seen a resurgence in recent years [7, 23, 13, 45], offering a means to estimate the effectiveness of diverse media channels using aggregated, observational data. This renewed interest is partly fueled by the increasing difficulty of long-term tracking for attribution and user-level experiments [32], especially given the deprecation of third-party cookies and stricter privacy regulations like GDPR, CCPA [6] and Apple’s Intelligent Tracking Prevention (ITP) [1]. MMMs circumvent these tracking needs by operating on aggregated data (e.g., weekly impressions and sales per region).
However, current MMM methods face well-documented limitations [7, 33]. First, MMM datasets are often limited in samples relative to the number of factors influencing sales, which can lead to statistically overdetermined models [7] which has historically limited the feasibility of employing more complex, data-hungry techniques like deep neural networks. Second, the common assumption [^1] There are a few exceptions such as [30]. of time-invariant response curves makes it challenging to account for variations in marketing effectiveness due to changes in ad creative quality, targeting strategies, or external factors like seasonality [^2] Such as changes in effectiveness during the holiday shopping season.. Finally, capturing long-term brand effects is notoriously difficult; parametric adstock functions typically used to model carryover effects often lose identifiability from the baseline beyond a few months.
To address these challenges, this paper introduces NNN (Next-Generation Neural Networks for Marketing Measurement), a novel paradigm leveraging recent advancements in deep learning and representation learning. This is an experimental approach which uses a different foundational modeling method compared to current MMM methodology. Some key differences include:
  • Rich Data Representation. Instead of relying on scalar inputs (e.g., spend), NNN utilizes high-dimensional embeddings to represent both marketing activities (e.g., Search Ads, YouTube Ads) and crucial organic signals (e.g., Google Search queries). These embeddings capture not only quantitative volume but also qualitative aspects like ad creative content or search query semantics, allowing the model to directly address the challenge of time-varying creative effectiveness.
@@
Our paper is structured as follows. We give an overview of related working influencing our approach (Related Work). Next, we discuss the data used (Datasets), and how they are formatted in our method (Data Structure). We give an overview of the computation of the model (Model Overview), then go into further details regarding the how traditional data is augmented and embedded (Input Data Representations). We discuss the intricacies of our Transformer (Transformer) and details about how the model makes the final predictions (Prediction Heads). We discuss training details (Model Training) and attribution methods (Sales Attribution). We give an overview of our experiments and then discuss the results and analysis (Experiments). Finally, we give a discussion about how our method compares to existing measurement approaches (Discussion), and give a concluding overview (Conclusion).
@@ 3 Related Work
Our proposed NNN framework builds upon and diverges from several strands of research in marketing measurement, primarily Media Mix Modeling, the application of neural networks in this domain, and the use of embeddings to represent media signals. This section situates NNN within this existing literature.
@@ 3.1 Media Mix Modeling Evolution
The field of Media Mix Modeling (MMM) provides the foundational context for NNN, offering established methods for assessing advertising ROI using historical, aggregate data [19, 4]. Recognizing limitations in early linear models, Bayesian Media Mix Modeling (BMMM) emerged, introducing techniques to model ad carryover and saturation (shape effects) more flexibly, often using Hill and adstock functions [23]. BMMM leverages Bayesian inference (e.g., MCMC methods) to incorporate prior knowledge into model estimation, a crucial aspect given the often data-constrained nature of MMM problems. This Bayesian framework was further extended to handle geographical variations through hierarchical modeling [34]. Meridian, a Google Bayesian MMM framework integrated advancements of MMM research over the past decade. It incorporates features like ROI priors and calibration [44], reach and frequency data integration [43], and paid search modeling. While these approaches are powerful and innovative, here we are exploring an entirely new research direction that does not rely on scalar inputs or pre-defined functional forms for temporal effects.
@@ 3.2 Neural Networks in Media Mix Modeling
The potential for neural networks (NNs) in MMM has been recognized for some time [36], owing to their ability to capture complex non-linearities. However, concerns over interpretability and the data requirements for training large NNs have historically limited their adoption in practice. Recent efforts aim to bridge this gap. For instance, CausalMMM [15] attempts to learn causal structures using Granger causality [18] and employs a variational autoencoder (VAE) [12] with a dedicated causal relational encoder and a marketing response decoder where the temporal response is learned alongside the saturation response. While innovative in its use of NNs and causality, this approach still operated on scalar representations of media activity. NNN builds on the potential of NNs but focuses on leveraging richer input representations via embeddings.
@@ 3.3 Embedding-Based Media Representations
A core innovation of NNN is its use of high-dimensional embeddings, drawing heavily on the work presented in [28]. That work demonstrated that organic signals, specifically aggregated Google Search query embeddings, could be highly predictive of real-world outcomes like flu rates and auto sales. This contrasted with earlier studies that used Google Search data but often relied on counts of specific keywords, necessitating significant feature engineering [9, 26, 25, 38, 40]. The key insight from [28] was using pre-trained language models to embed search terms and then aggregating these embeddings (via summation in SLaM) to create a dense “search embedding” for a given time and geo, capturing both the volume and semantic content of search activity as a proxy for organic consumer demand. In our work, we show that the search embedding is a highly effective baseline for regression-based measurement models, and we introduce additional embeddings for other paid marketing channels such as Google Search ads and YouTube ads.
@@ 3.4 Terminology and Conventions
Throughout this paper, we adopt specific terminology for clarity. We use “MMM” to refer generally to current Marketing Mix Modeling methods, which could be Bayesian or Frequentist, and generally use impressions or other scalar quantities to represent media. In contrast, “NNN” denotes our proposed neural network-based methodology which uses embeddings and scalars to represent media.
@@ 4 Datasets
    Data was randomly generated using the test_utils module of the Meridian package to simulate 130 weeks of data over 100 random geos. We used this module for simulated data generation given its ease of use, but no properties of the simulated data are specific to the package. We created two channels of random media data (YouTube and Search Ads) and one channel of random controls data (Search), applying adstock and Hill functions to the media channels. In order to model sales driven directly by YouTube, we assumed an adstock retention rate $\alpha_{m}=0.75$ and a Hill function effective concentration $e_{c}=1.0$ and applied these functions to the random YouTube media data with arbitrary normalization. For the indirect sales driven by YouTube, we applied a different set of adstock and Hill functions to the YouTube media data ($\alpha_{m}=0.5$ and $e_{c}=3$) and added those normalized impressions to the random search data so that they contributed an additional 20% or so on top of the random search data. We then generated sales from the sum of the random search impressions and the YouTube-generated search impressions, assuming a flat conversion rate of 0.1%. Finally, search ads (which in this simplified model are not causally connected to search) also generated sales using $\alpha_{m}=0.3$ and $e_{c}=1$.
@@
    In the end we have two versions of this dataset: High variance and low variance, differing only in the magnitude of the random noise corresponding to “intent” or creative effectiveness. For the high-variance dataset, this varies uniformly between 0.5 and 1.5, while in the low-variance dataset it varies uniformly between 0.8 and 1.2. We compare our method against a baseline built using current MMM methods on these datasets on both sales-prediction error-rate evaluation and on sales channel attribution estimates, because the ground-truth attributions are known when creating the data.
  • Real-world, Regional-level. (4 regions). This data is real-world marketing and sales data, aggregated at a regional- and weekly-level. We use it for an apples-to-apples comparison of NNN against a baseline MMM in a real-world setting where priors have already been established. We use only the marketing volume as model inputs (quantitative information) in both NNN and the baseline MMM. The Search embeddings are generated using approximately 8M queries that are categorically related to the business’s sales.
@@ 6 Model Overview
Our baseline NNN model is based on the Search-to-Sales dependency in Figure 1. Previous work has shown that Google Search is a good proxy for the consumer demand, and even captures seasonal effects; we use an embedded version of Search where user queries are aggregated according to their semantic representations from a language model [28]. We pass this embedded representation of search to a function that is similar to the hill functions found in many modern MMMs to get our estimate of Sales. We enrich our Sales estimate by including marketing-to-search effects (the blue lines from Marketing to Sales in 1) in a similar fashion, utilizing embedded versions of the paid media. To include the temporal effects in Figure 1 (the yellow lines from Search to Sales and the blue lines from Marketing to Sales), we utilize a Transformer architecture that is able to attend to channels’ histories. This allows the feature representation used to predict sales to depend not just on the current time-period’s Search, but also it’s history. Similarly, this is done for all other channels. Finally, our model is allowed to be trained not just to predict sales, but to predict organic channels like Search (the grey lines in Figure 1). This allows our model to capture marketing effects that happen through an intermediary such as Google Search (E.g., Imagine a channel that affects Search, and which then drives sales).
@@ 7.5 Scalars
In the case where there is no way to use a model to embed data (e.g., in the scenario where only impressions of a channel are stored, and no other contextual information is stored along side it, perhaps for privacy reasons or legacy logging methods), similar to MMM we use only the scalar quantity in our models. Our architecture requires all channels to be represented as a $D$-dimensional value, so we pad the scalar with $D-1$ zero values for the other dimensions. Note that the magnitude of this vector is still the channel volume, which is consistent with all other channels. In this paper we demonstrate one example of independent media variables modeled this way (the real-world dataset), and of course sales (the dependent variable) is also a scalar in our models.
@@ 8.1 Attention
In our model, we use historical information across marketing channels to predict sales or other business KPIs on a given day $t$. A key innovation in our NNN framework is that we do not have to specify a functional form for the time lag between media exposure and sales. Instead, we employ attention mechanisms [2, 39] that learn which channels and prior time steps are most relevant for the current prediction context (channel $c$, time $t$).
@@ 9.1 Sales Head
and the function $P$ is an MLP Resnet where the final layer has one hidden unit and a sigmoid activation function with parameters $\theta^{\text{Search}}$. The effect of this roughly mimics a curve of diminishing returns (often modeled by MMMs using a Hill or Weibull function) in that the marginal effectiveness of the impressions on sales saturates. A key difference, however, is that in MMMs the saturation is with respect to the advertising volume. In this model, volume scales linearly without saturation, and the output of the saturated “intent” function represents the conversion rate of a Search query.
To add multi-channel flexibility in our model, we incorporate the effect of more marketing channels additively [^5] We experimented with using multiplicative effects via using a log-scale version of the model, but we found the additive models generally performed better and match our expectations that most marketing channels are independent of one another; see the Appendix for more details as
@@ 9.2 Search Head
It is generally understood that intermediary organic channels such as Google Search are affected via marketing, but these are often not modeled explicitly. We aim to incorporate these effects into NNN by generalizing our modeling framework to predict not just sales, but also the Search information. We do so by predicting the search embedding $X_{g,t+1,\text{Search}}$ at time step $t$ for geo $g$, learning both the magnitude and direction of the search [^7] This means our model output has values that are representative of the following time-step for the Search channel.. This allows us to capture movements in search patterns due to marketing (i.e., movements in the “baseline” sales) which are generally treated as exogeneous variables in current MMMs. Our search head looks like
@@ 9.3 Other Heads
The other inputs to our model are marketing channels, which are controlled inputs, whereas the Search and Sales channels are measurements / observations. Because we treat these as marketing levers and not observations (although it is a viewpoint that Search can affect Search ads, we neglect this in our current models), we don’t have a need to predict their values during training. Out of convenience, to keep the tensor dimensions consistent, we project the intermediate representations of all other channels back to their original size.
@@ 10 Model Training
We used open-source Bayesian-based MMM methods as a baseline point of comparison, using default out-of-the-box parameters. For the real-world experiments we use priors from a previous training of the model, which replicates a common practice with real-world MMMs.
@@ 11 Sales Attribution
where $X$ represents the original data and media volume and $\tilde{X}^{c}$ represents the same data except the media from channel $c$ is set to zero. Because there can be bias errors in predictions, we report the predictions on a mix or percentage basis, defined as $\text{mix}\%_{c}=\frac{\text{Attribution}_{c}}{\sum_{c^{\prime}}\text{Attribution}_{c}^{\prime}}$.
@@
Since our framework also allows us to model the effect on the intermediary organic channels, we can autoregressively unroll predictions out in time. We start with a prefix $P$ of data and compute the following